OpenStudy (anonymous):

Identify square root of 2 as either rational or irrational, and approximate to the tenths place.

OpenStudy (anonymous):

i got it

OpenStudy (zzr0ck3r):

The square root of a prime number is irrational. Go to wolframalpha.com and type sqrt{2} for an approximation

OpenStudy (anonymous):

irrational 1.4

OpenStudy (zzr0ck3r):

correct

OpenStudy (anonymous):

can i ask u another question?

OpenStudy (zzr0ck3r):

sure

OpenStudy (anonymous):

Solve x3 = 1 over 8. 1 over 2 ±1 over 2 1 over 4 ±1 over 4

OpenStudy (zzr0ck3r):

\(x^3=\dfrac{1}{8}\implies x=\sqrt[3]{\dfrac{1}{8}}\implies x=\dfrac{\sqrt[3]{1}}{\sqrt[3]{8}}\implies x=\dfrac{1}{\sqrt[3]{8}}= \ ?\)

OpenStudy (anonymous):

im confuzed

OpenStudy (zzr0ck3r):

Which step?

OpenStudy (anonymous):

2nd one

OpenStudy (zzr0ck3r):

1) I took the cube root of both sides which gives just \(x\) on the left and the cube root of the fraction on the right 2) the cube root of a fraction is the cube root of the numerator over the cube root of the denominator. 3) the cube root of \(1\) is \(1\). 4) what is the cube root of \(8\)?

OpenStudy (anonymous):

2.82

OpenStudy (anonymous):

is it C?

OpenStudy (zzr0ck3r):

what number times itself 3 times equals 8?

OpenStudy (anonymous):

2

OpenStudy (zzr0ck3r):

so \(x=\dfrac{1}{2}\)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

WaiT!!!!! @zzr0ck3r

OpenStudy (anonymous):

is it C or D im confussled

OpenStudy (anonymous):

i mean A OR B